# EDINET-Bench: Evaluating LLMs on Complex Financial Tasks using Japanese Financial Statements

### v1

@@ Abstract
Financial analysis presents complex challenges that could leverage large language model (LLM) capabilities. However, the scarcity of challenging financial datasets, particularly for Japanese financial data, impedes academic innovation in financial analytics. As LLMs advance, this lack of accessible research resources increasingly hinders their development and evaluation in this specialized domain. To address this gap, we introduce EDINET-Bench, an open-source Japanese financial benchmark designed to evaluate the performance of LLMs on challenging financial tasks including accounting fraud detection, earnings forecasting, and industry prediction. EDINET-Bench is constructed by downloading annual reports from the past 10 years from Japan’s Electronic Disclosure for Investors’ NETwork (EDINET) and automatically assigning labels corresponding to each evaluation task. Our experiments reveal that even state-of-the-art LLMs struggle, performing only slightly better than logistic regression in binary classification for fraud detection and earnings forecasting. These results highlight significant challenges in applying LLMs to real-world financial applications and underscore the need for domain-specific adaptation. Our dataset, benchmark construction code, and evaluation code is publicly available to facilitate future research in finance with LLMs.
Figure 1: Overview of EDINET-Bench.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable capabilities across various domains, including general knowledge, mathematics, and coding (OpenAI, 2024; Anthropic, 2024; Google, 2024; DeepSeek-AI, 2025). LLMs are also being explored for applications in the financial sector, where they have shown strong performance in various benchmark tasks (Wu et al., 2023; Yang et al., 2023; Xie et al., 2023). However, most existing benchmarks focus on relatively simple tasks such as information extraction and question answering (Chen et al., 2021, 2022b; Islam et al., 2023).
To make a significant societal impact in the financial domain, LLMs must excel in expert-level decision-making tasks such as accounting fraud detection and earnings forecasting (Beneish, 1999; Song et al., 2016; Kondo et al., 2019; Penman and Sougiannis, 1998; Monahan, 2018). In addition, current financial benchmarks predominantly focus on English-speaking markets (US and Europe) or Chinese-language content. Japan, the fourth largest economy globally by nominal GDP (in 2023; World Bank (2025)), plays a crucial role in the global economy, yet remains underserved and underrepresented by AI technology developments. Japanese financial benchmarks could significantly advance digital transformation efforts throughout the country.
To address these gaps, we introduce EDINET-Bench, an open-source Japanese financial benchmark designed to evaluate LLMs on complex financial tasks. We construct EDINET-Bench using Japan’s Electronic Disclosure for Investors’ NETwork (EDINET) [^1] https://disclosure2.edinet-fsa.go.jp/, an electronic system that digitizes and publicly discloses securities reports to improve market efficiency and information accessibility. EDINET-Bench features challenging tasks including accounting fraud detection and earnings forecasting, which require sophisticated financial understanding. Our benchmark is distinctive for its automated construction and exclusive Japanese language content. As the dataset construction is automatic, we can easily incorporate future annual reports.
Using EDINET-Bench, we evaluate the performance of frontier LLMs in a zero-shot setting. Our results indicate that even state-of-the-art models struggle, performing only slightly better than logistic regression in binary classification for fraud detection and earnings forecasting. These results highlight current LLM limitations in complex financial analysis and underscore the need for dedicated research to address these challenges. By open-sourcing our benchmark dataset and toolkit, we enable researchers to tackle challenging financial tasks using AI and facilitate continuous development of the benchmark.
@@ 2 Background and related work
Financial benchmarks
Many financial benchmarks have been proposed in order to evaluate the financial knowledge and reasoning abilities of LLMs (Chen et al., 2021; Islam et al., 2023; Xie et al., 2024; Guo et al., 2025). For example, FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022b) are closed numerical reasoning question answering tasks related to financial analysis. Similarly, FinanceBench (Islam et al., 2023) is an open-book financial question answering benchmark dataset, and FinBen (Xie et al., 2024) is a benchmark-suit that contains 24 tasks. FAMMA (Xue et al., 2025) uses university textbooks and the CFA exam to construct a multimodal question answering benchmark. Several benchmarks exist in the Japanese financial domain (see Nakagawa et al. (2025) for a recent survey), including Japanese-lm-fin-harness (Hirano, 2024) and JCPA dataset (Masuda et al., 2023), which focus on Japanese financial exam questions. Most of these datasets rely on financial experts to design high-quality questions, and therefore cannot be easily extended to include up-to-date questions after the publication. On the other hand, our framework can easily incorporate future financial documents, and automatically annotates the labels for each task.
Accounting fraud detection
@@
Accounting fraud detection involves identifying deliberate misstatements, manipulations, or discrepancies in financial reports. These reports are critical for numerous stakeholders, as investors base allocation decisions on them, while job seekers evaluate potential employers through their financial disclosures (Sou, 2018). Despite preliminary audits, fraudulent activities are frequently discovered only after publication, necessitating subsequent corrections (Japan Institute of Certified Public Accountants, 2024). Accounting fraud detection has been studied for many years, with pioneering work by Beneish (1999). After that, traditional machine learning approaches have been applied to accounting fraud detection (Perols, 2011; Dechow et al., 2011; Song et al., 2016; West and Bhattacharya, 2016; Kondo et al., 2019), but effectively processing textual information alongside numerical data in annual reports using LLMs remains underexplored.
Earnings forecasting
@@
While earnings forecasting typically involves predicting the numeric magnitude of future earnings (Penman and Sougiannis, 1998; Monahan, 2018), a large body of work instead focuses on predicting the sign of the change in earnings (Ou and Penman, 1989; Chen et al., 2022a), which remains a difficult task for professionals. Kim et al. (2024) investigated if GPT-4, when provided with financial statements in which company names and fiscal years are anonymized, and without using any textual information, will outperform human analysts in predicting the direction of earnings for the following year. However, it is based on proprietary data. On the other hand, we will publish both the evaluation dataset and evaluation code to facilitate future research in this area.
Industry prediction
@@
Industry prediction is a multi-class classification task that aims to predict the industry category based on security reports. While listed firms already are tagged with industry labels, e.g., by Securities Identification Code Committee (SICC), portfolio managers may want to adopt their own industry definitions, e.g., see Kimura and Nakagawa (2022) for a data-driven approach, or anticipate reclassifications as firms’ business evolve or mergers reshape their operations. Evaluating whether LLMs can predict industries based on financial information such as balance sheets and profit/loss statements serves as an effective task for measuring LLMs’ financial domain knowledge. Van Der Heijden (2022); Dolphin et al. (2023) used machine learning for this application.
Table 1: Sample of TSV records. Some columns are omitted for ease of reading.
Element Name Relative Year Unit Value
提出回数 (Number of Submissions) 提出日時点 (As of Submission Date) 1
売上高 (Sales) 前期 (Previous year) 円 (Yen) 10000000
売上高 (Sales) 今期 (Current year) 円 (Yen) 20000000
EDINET
@@
Electronic Disclosure for Investors’ NETwork (EDINET) is a platform managed by the Financial Services Agency (FSA) of Japan that provides access to disclosure documents such as securities reports. This platform offers both a web interface and an EDINET API to access reports, allowing users to download annual, semi-annual, quarterly, and amended reports for the past ten years. In addition to PDFs, TSV files are also available, containing structured data where each row represents a record of individual attributes that exist in the annual report PDF. Table 1 shows examples of TSV records. Detailed information for each disclosure item can be extracted and analyzed by parsing these attributes. EDINET operates as the equivalent of EDGAR [^2] https://www.sec.gov/edgar/search/ for the United States. We use EDINET as the primary data source for our benchmark.
@@ 3 Construction of EDINET-Bench
In this section, we describe the construction method of EDINET-Bench. Overview of our construction pipeline is shown in Figure 1. Generally, EDINET-Bench is constructed by downloading Japanese listed companies’ annual reports from EDINET and automatically assigning labels for each task.
@@
We first create edinet2dataset, a tool for downloading and parsing financial documents from EDINET. This tool includes a download function for securities reports using the EDINET API and a parsing function to extract financial information from the downloaded report files in TSV format. The parsing process leverages Polars [^3] https://github.com/pola-rs/polars to enable high-speed processing. edinet2dataset is inspired by the edgar-crawler [^4] https://github.com/lefterisloukas/edgar-crawler from the Edgar-Corpus (Loukas et al., 2021) and shares similar functionality. However, edinet2dataset is capable of extracting more detailed information, such as the current year’s sales. The information extracted by edinet2dataset can be broadly categorized as follows: Meta: Metadata such as company name. Summary: Key financial indicators. BS: Balance sheet statement. PL: Profit and loss statement. CF: Cash flow statement. Text: Other textual information in the report, such as company history, business risk explanations, and auditor information. Appendix B shows an example of parsed items from an annual report using edinet2dataset.
@@ 3.2 EDINET-Corpus
By using edinet2dataset, we collect all available annual reports and amended annual reports from EDINET for the period from April 2014 to April 2025, which includes approximately 40,000 documents (4,000 listed companies $\times$× 10 years). The number of annual reports obtained from EDINET over the past ten years for each year, as well as the number of companies per industry, are shown in Table 2 and Figure 2. We call this dataset the EDINET-Corpus. The EDINET-Corpus serves as the primary data source for constructing EDINET-Bench, as described in the next section.
@@ 3.3 EDINET-Bench
Table 5: Class distribution in industry prediction task.
@@
Accounting fraud detection
We construct a binary classification dataset for accounting fraud detection. We collect both fraudulent and non-fraudulent reports in order to build this dataset. For the fraudulent reports, we first downloaded the past ten years of amended annual reports from EDINET, obtaining a total of 6,712 reports. An amended annual report is a document that discloses amendments to the original annual report. These amendments can be related to fraudulent activities or to non-fraudulent issues, such as misreporting the ownership percentage and share count of major shareholders, under-reporting executive compensation, or failing to disclose critical information. To identify amendments linked to fraudulent activities, we extract text from amended annual reports in PDF format using pdfminer [^5] https://pypi.org/project/pdfminer/, a tool for PDF text extraction. We then analyze the extracted text using Claude 3.7 Sonnet, prompting the model to determine whether amendments are associated with fraud. The specific prompt used for this classification is detailed in Appendix C. As a result, 668 reports were identified as fraudulent and are labeled as fraudulent. Note that our prompt leaves room for unintentional errors, such as misstatements, to be included. According to our manual check, our fraudulent set consists mostly of intentional manipulations, but a small proportion of the reports are unintentional errors. For the non-fraudulent reports, we randomly sample 700 companies from all annual reports submitted over the past ten years, excluding those identified as fraudulent. From each selected company, we randomly choose one annual report from the reports of the company that have been downloaded. We then split the fraudulent and non-fraudulent samples into training and test sets, ensuring that data from the same company does not overlap between the two sets. This resulted in a dataset with 865 samples for training and 224 samples for testing, totaling 1,089 instances. Due to document parsing errors, the final dataset was reduced to 534 fraudulent and 555 non-fraudulent samples, smaller than the initial 668 and 700 samples, respectively. The class distribution of the final dataset is shown in Table 5.
Earnings forecasting
@@
Industry prediction
We construct a multi-class classification dataset for industry prediction to predict a company’s industry based on its annual report. We use the 33-class industry information provided by EDINET for each company, also known as TOPIX-33 and is widely recognized among Japanese investors. The industry definitions follow the classification determined by the SICC, which is based on the Japan Standard Industrial Classification published by the Ministry of Internal Affairs and Communications (MIC). However, the granularity of 33 industries is considered too fine for our purposes. To enhance interpretability and simplify the prediction task, we consolidate similar industries into 16 broader categories [^8] In TOPIX-17, the number of cases in the industries of “Energy Resources” and “Electricity & Gas,” which are loosely related domains, was smaller compared to other industries. Therefore, these two were merged to create 16 industries.. We categorize all companies available in the EDINET-Corpus into 16 industry groups and randomly sample approximately 35 companies from each category. For each sampled company, we use reports with fiscal year starting from 2023. This process yielded a dataset of 496 samples with the class distribution shown in Table 5. In this task, since the minimum number of companies per industry among the 16 industries was only 23, which is relatively small, we did not split the data into training and test sets.

### v2

@@ Abstract
Large Language Models (LLMs) have made remarkable progress, surpassing human performance on several benchmarks in domains such as mathematics and coding. A key driver of this progress has been the development of benchmark datasets. In contrast, the financial domain poses higher entry barriers due to its demand for specialized expertise, and benchmarks remain relatively scarce compared to those in mathematics or coding. We introduce EDINET-Bench, an open-source Japanese financial benchmark designed to evaluate LLMs on challenging tasks such as accounting fraud detection, earnings forecasting, and industry classification. EDINET-Bench is constructed from ten years of annual reports filed by Japanese companies. These tasks require models to process entire annual reports and integrate information across multiple tables and textual sections, demanding expert-level reasoning that is challenging even for human professionals. Our experiments show that even state-of-the-art LLMs struggle in this domain, performing only marginally better than logistic regression in binary classification tasks such as fraud detection and earnings forecasting. Our results show that simply providing reports to LLMs in a straightforward setting is not enough. This highlights the need for benchmark frameworks that better reflect the environments in which financial professionals operate, with richer scaffolding such as realistic simulations and task-specific reasoning support to enable more effective problem solving. We make our dataset and code publicly available to support future research.
Figure 1: EDINET-Bench is a challenging benchmark evaluating LLMs on fraud detection, earnings forecasting, and industry classification from annual reports with text and tables.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable capabilities across diverse domains, including general knowledge, mathematics, and coding (achiam2023gpt; claude35sonnet; gemini1_5; liu2024deepseek; kimiteam2025kimik2). A key driver of this progress is the development of benchmark datasets. By evaluating models on benchmarks, researchers can identify current limitations, refine model architectures, training data, and methods, and ultimately climb the “benchmark hill”. As a result, many existing benchmarks are now approaching saturation (OpenAI2025), motivating the creation of new benchmarks that are substantially more difficult and less prone to rapid saturation.
Prominent examples include Humanity’s Last Exam (phan2025humanitysexam), which consists of expert-level QA problems; SWE-Bench variants (jimenez2024swebenchlanguagemodelsresolve; chowdhury2024swebenchverified; deng2025swebenchproaiagents), which require fixing real GitHub issues so that unit tests pass; and WebArena (zhou2024webarenarealisticwebenvironment), which evaluates agents operating real web browsers. Such benchmarks push models closer to real-world usability by mirroring expert-level tasks.
While benchmark research has advanced across many domains, progress in finance has lagged behind (Lee2025surveyfinllm). Existing financial benchmarks primarily focus on relatively simple tasks, such as knowledge-based QA or data extraction from tables (chen2021finqa; chen2022convfinqa; islam2023arxiv). These are valuable for measuring basic domain knowledge but fall short of the kind of complex, high-stakes tasks that, if solved, would immediately translate to real-world utility.
To fill this gap, we introduce EDINET-Bench, an open-source financial benchmark constructed from ten years of real-world financial reports filed through Japan’s Electronic Disclosure for Investors’ NETwork (EDINET) [^1] https://disclosure2.edinet-fsa.go.jp/ . EDINET-Bench presents demanding financial analysis challenges, including accounting fraud detection, earnings forecasting, and industry classification. These tasks require models to process complete annual reports, combine information from multiple tables and text sections, and demonstrate expert-level financial reasoning. To the best of our knowledge, this is the first open dataset and benchmark for accounting fraud detection.
We evaluate frontier LLMs on EDINET-Bench in zero-shot settings. Our results show that even state-of-the-art models struggle, performing only marginally better than logistic regression on binary classification tasks. This suggests that merely providing annual reports to LLMs in a simple setting is insufficient. Our findings underscore the need for benchmark frameworks that more closely mirror the environments in which financial professionals operate, incorporating richer scaffolding such as realistic simulations and task-specific reasoning support to facilitate more effective problem solving.
By releasing EDINET-Bench and its construction toolkit, we provide a foundation for studying expert-level financial reasoning with LLMs and for developing more powerful methods that can eventually support real-world financial applications.
Table 1: Comparison of financial benchmarks. Expert-level reasoning refers to tasks that require synthesizing information across tables and text, including analyzing year-over-year trends, interpreting relationships between multiple financial tables, and combining these signals with narrative disclosures. Tasks such as fraud detection and earnings forecasting fall into this category, as they require comprehensive processing of both tabular and textual evidence together with domain-specific financial and accounting expertise.
Benchmark Input modality Expert level reasoning Language
FinQA (chen2021finqa) Text, Table En
ConvFinQA (chen2022convfinqa) Text, Table En
FinanceBench (islam2023arxiv) Text En
FinanceMATH (zhao-etal-2024-knowledgefmath) Text, Table En
FAMMA (xue2024famma) Text, Image En
AuditBench (wang2024automating) Text, Table En
Japanese-lm-fin-harness (hirano2023finnlpkdf) Text, Table Ja
EDINET-Bench (Ours) Text, Table Ja
@@ 2 Background and related work
Financial benchmarks.
A variety of financial benchmarks have been proposed in order to evaluate the financial knowledge and reasoning abilities of LLMs (chen2021finqa; islam2023arxiv; xie2024neuripsdb; guo2024finevalchinesefinancialdomain). For example, FinQA (chen2021finqa) and ConvFinQA (chen2022convfinqa) are closed numerical reasoning question answering tasks related to financial analysis. Similarly, FinanceBench (islam2023arxiv) is an open-book financial question answering benchmark dataset, and FinBen (xie2024neuripsdb) is a benchmark-suit that contains 24 tasks. FAMMA (xue2024famma) uses university textbooks and the CFA exam to construct a multimodal question answering benchmark. Several benchmarks exist in the Japanese financial domain (see nakagawa2025survey for a recent survey), including Japanese-lm-fin-harness (hirano2023finnlpkdf) and JCPA dataset (masuda2023jcpa), which focus on Japanese financial exam questions. Recently, benchmarks such as FinanceMATH (zhao-etal-2024-knowledgefmath), which evaluates an LLM’s financial knowledge together with its mathematical reasoning, and AuditBench (wang2024automating), which identifies subtle discrepancies between transaction data and financial statements, have been introduced. Despite these advances, existing benchmarks still center on tasks that can be solved with basic financial knowledge and relatively simple computations. As summarized in Table 1, most existing benchmarks rely on relatively simple QA settings that do not require expert-level reasoning. In contrast, EDINET-Bench poses a more complex challenge: the inputs consist of entire financial reports, and solving tasks such as accounting fraud detection requires integrating information across multiple tables and textual sections, demanding expert-level reasoning.
Accounting fraud detection.
@@
Accounting fraud detection involves identifying deliberate misstatements, manipulations, or discrepancies in financial reports. These reports are critical for numerous stakeholders, as investors base allocation decisions on them, while job seekers evaluate potential employers through their financial disclosures (sou2018fsa). Despite preliminary audits, fraudulent activities are frequently discovered only after publication, necessitating subsequent corrections (jicpa2024fraud). Accounting fraud detection has been studied for many years, with pioneering work by beneish1999detection. After that, traditional machine learning approaches have been applied to accounting fraud detection (perols2011financial; dechow2011predicting; song2016predicting; WEST201647; kondo2019rieti), but effectively processing textual information alongside numerical data in annual reports using LLMs remains underexplored.
Earnings forecasting.
@@
While earnings forecasting typically involves predicting the numeric magnitude of future earnings (penman1998comparison; manahan2018earnings), a large body of work instead focuses on predicting the sign of the change in earnings (ou1989financial; chen2022predicting), which remains a difficult task for professionals. kim2024financialstatementanalysislarge investigated if GPT-4, when provided with financial statements in which company names and fiscal years are anonymized, and without using any textual information, will outperform human analysts in predicting the direction of earnings for the following year. However, it is based on proprietary data and evaluation code were not made public. On the other hand, we will publish both the evaluation dataset and evaluation code to facilitate future research in this area.
Industry prediction.
@@
Industry prediction is a multi-class classification task that aims to predict the industry category based on security reports. While listed firms already are tagged with industry labels, e.g., by Securities Identification Code Committee (SICC), portfolio managers may want to adopt their own industry definitions, e.g., see kimura2022industry for a data-driven approach, or anticipate reclassifications as firms’ business evolve or mergers reshape their operations. Evaluating whether LLMs can predict industries based on financial information such as balance sheets and profit/loss statements serves as an effective task for measuring LLMs’ financial domain knowledge. van2022predicting; dolphin2023aics used machine learning for this application.
EDINET.
@@
Electronic Disclosure for Investors’ NETwork (EDINET) is a platform managed by the Financial Services Agency (FSA) of Japan that provides access to disclosure documents such as securities reports. This platform offers both a web interface and an EDINET API to access reports, allowing users to download annual, semi-annual, quarterly, and amended reports for the past ten years. In addition to PDFs, TSV files are also available, containing structured data where each row represents a record of individual attributes that exist in the annual report PDF. Detailed information for each disclosure item can be extracted and analyzed by parsing these attributes. EDINET is similar to EDGAR [^2] https://www.sec.gov/edgar/search/ in the United States. We use EDINET as the primary data source for our benchmark.
@@ 3 Construction of EDINET-Bench
In this section, we describe the construction method of EDINET-Bench. Overview of our construction pipeline is shown in Figure 1. Generally, EDINET-Bench is constructed by downloading Japanese listed companies’ annual reports from EDINET and assigning labels for each task.
@@
We first create edinet2dataset, a tool for downloading and parsing financial documents from EDINET. This tool primarily includes a download function for securities reports using the EDINET API and a parsing function to extract financial information from the downloaded report files in TSV format. The parsing process leverages Polars (polars) to enable high-speed processing. edinet2dataset is inspired by the edgar-crawler from the Edger-Corpus (loukas-etal-2021-edgar) and shares similar functionality. However, edinet2dataset is capable of extracting more detailed information, such as the current year’s sales. The information extracted by edinet2dataset can be broadly categorized as follows: Meta: Metadata such as company name and EDINET code. Summary: Key financial indicators. BS: Balance sheet statement. PL: Profit and loss statement. CF: Cash flow statement. Text: Other textual information in the report, such as company history and business risk explanations.
@@ 3.2 EDINET-Corpus
By using edinet2dataset, we collect all available annual reports and amended annual reports from EDINET for the period from April 2014 to April 2025, which includes approximately 40,000 documents (4,000 listed companies $\times$ 10 years). The number of annual reports obtained from EDINET over the past ten years for each year, as well as the number of companies per industry, are shown in Table 2 and Figure 2. We call this dataset the EDINET-Corpus, and it will be publicly available for reproducibility. The EDINET-Corpus serves as the primary data source for constructing EDINET-Bench, as described in the next section.
@@ 3.3 EDINET-Bench
Table 5: Class distribution in industry prediction task (test split).
@@
Accounting fraud detection.
We construct a binary classification dataset for accounting fraud detection. We collect both fraudulent and non-fraudulent reports in order to build this dataset. For the fraudulent reports, we first downloaded the past ten years of amended annual reports from EDINET, obtaining a total of 6,712 reports. An amended annual report is a document that discloses amendments to the original annual report. These amendments can be related to fraudulent activities or to non-fraudulent issues, such as misreporting the ownership percentage and share count of major shareholders, under-reporting executive compensation, or failing to disclose critical information. To identify amendments linked to fraudulent activities, we extract text from amended annual reports in PDF format using pdfminer (pdfminer) , a specialized tool for PDF text extraction. Notably, amended reports explicitly state the reasons for corrections in text form, which provides direct evidence of irregularities. Leveraging this property, we analyze the extracted text with Claude 3.7 Sonnet, prompting the model to assess whether the stated reasons indicate fraud. This enables reliable and large-scale detection of fraudulent cases. The specific prompt used for this classification is detailed in Appendix B. As a result, 668 reports were identified as fraudulent and are labeled as fraudulent. We manually reviewed all cases labeled as fraudulent by reading the amendment reasons in the reports. Fewer than 5% of the fraud-labeled cases appeared unrelated (e.g., board member changes), suggesting a low error rate and supporting the reliability of the labels. For the non-fraudulent reports, we randomly sample 700 companies from all annual reports submitted over the past ten years, excluding those identified as fraudulent. From each selected company, we randomly choose one annual report from the reports of the company that have been downloaded. We then split the fraudulent and non-fraudulent samples into training and test sets, ensuring that data from the same company does not overlap between the two sets. This resulted in a dataset with 865 samples for training and 224 samples for testing, totaling 1,089 instances. Due to document parsing errors, the dataset was reduced to 534 fraudulent and 555 non-fraudulent samples, smaller than the initial 668 and 700 samples, respectively. The class distribution of the dataset is shown in Table 5.
Earnings forecasting.
@@
Industry prediction.
We construct a multi-class classification dataset for industry prediction to predict a company’s industry based on its annual report. We use the industry information provided by EDINET for each company. The industry definitions follow the classification established by the SICC, which categorizes listed companies into 33 distinct industries. This classification, known as TOPIX-33, is widely recognized among Japanese investors. However, the granularity of 33 industries is considered too fine for our purposes. To enhance interpretability and simplify the prediction task, we consolidate similar industries into 16 broader categories [^5] In TOPIX-17, the number of cases in the industries of “Energy Resources” and “Electricity & Gas,” which are loosely related domains, was smaller compared to other industries. Therefore, these two were merged to create 16 industries.. This merging process is guided by the Japan Standard Industrial Classification published by the Ministry of Internal Affairs and Communications (MIC) and the SICC’s official guidelines. We categorize all companies available in the EDINET-Corpus into 16 industry groups and randomly sample approximately 35 companies from each category. For each sampled company, the latest fiscal year’s securities report is used as input for the industry prediction task. This process yields a test split of 496 samples with the class distribution shown in Table 5. In addition, we prepare a train split consisting of 900 examples.